1. A comet with a mass of 75,000 kg is moving with a velocity of 25,000 m/s. Calculate its kinetic energy.
1
Expert's answer
2021-02-02T07:07:01-0500

Solution:

Equation of kinetic energy is

"KE=\dfrac{mv^2}{2}"



Given:

m = 75,000 kg

v = 25,000 "(\frac{m}{s})"



"KE=\dfrac{(75,000kg)(25,000(\frac{m}{s}))^2}{2}"



Answer:

The comet’s kinetic energy is "KE = 23437.5 J"